The metal potassium crystallizes in a body centered cubic unit cell with one atom per lattice point. When X-rays with λ = 1.937 Å are used, the second-order Bragg reflection from a set of parallel planes in a(n) potassium crystal is observed at an angle θ = 21.32°. If the spacing between these planes corresponds to the unit cell length (d = a), calculate the radius of a(n) potassium atom.
Answer #1

From Bragg's Law,

n\lambda = 2d sin\theta

Where n = order of Bragg's reflection = 2

\theta = angle = 21.32o

\lambda =1.937 \AA

So, 2*(1.937 \AA) = 2*d*sin(21.32)

d = 5.328\AA = a= unit cell length

From BCC, relation between a and radius of K atom are:

d = 4*r / \sqrt{3}

5.328\AA = 4*r / \sqrt{3}

r = 2.307\AA

Hence, radius of a potassium atom = 2.307\AA